a formula used to compute the current value of a savings account is a= p(l+r)^n, where a is the current value; p is the amount deposited; r is the rate of interest for 1 compounding period, expressed as a decimal; and n is the number of compounding periods. which of the following is closest to the value of a savings account after 5 years if $10,000 is deposited at 4% annual intrest compounded yearly?
f. $10,400
g. $12,167
h. $42,000
j. $52,000
k. $53,782

Answer by rfer(16322)

A=10000(1+0.04)^5
A=10000(1.04)^5
A=10000(1.2166529)
A=$12,167